OpenStudy (anonymous):

What is the value of the binary number 110 in base 10?

OpenStudy (raden):

110 (2) = 1 * 2^2 + 1 * 2 + 0 * 2^0 = ...

OpenStudy (anonymous):

\(\large \color{red}{110}_2=\color{red}1*2^2+\color{red}1*2^1+\color{red}0*2^0=... \)

OpenStudy (anonymous):

6?

OpenStudy (acxbox22):